OpenStudy (anonymous):

Solve for x: 3/4 x +5/8= 4x x = __________________ Write your answer as a fraction in simplest form. Use the "/" symbol for the fraction bar.

OpenStudy (anonymous):

so x=20/125 and x=4/25 both the answers?

OpenStudy (anonymous):

or just x=4/25

OpenStudy (jacob902):

the answers x=5/26

OpenStudy (anonymous):

thnx :)

OpenStudy (anonymous):

wait 5/26 or 4/25?

OpenStudy (jacob902):

(3/4x)+(5/8)=4x And that is what I will solve. Multiply everything by 8: 6X+5=32X Move the 6X to the other side. When it's on the left, its sign is + when it moves, its sign becomes - +5=32X-6X Solve: 5=26X X=5/25

OpenStudy (jacob902):

5/26

OpenStudy (jacob902):

is the answer

OpenStudy (anonymous):

oh ok thnx :)